Currently you have $60 and your sister has $135. You decide to save $5 of your allowance each week, while your sister decides to spend her whole allowance plus $10 each week. How long will it be before you have as much money as your sister?
The answer is 5 weeks (I think) but I don't know how to set the equation up. Thanks.

Answer by gonzo(654)

you're right.
answer is 5 weeks.
here's what i think.
you start off with 60 and you add 5 each week.
she starts off with 135 and she subtracts 10 each week.
in x weeks you and her will be even.
so, your 60 plus 5 * x weeks will equal her 135 - 10 * x weeks.
set it up as follows:
-----
60 + 5*x = 135 - 10*x
add 10*x to both sides and subtract 60 from both sides and equation becomes
15*x = 135 - 60 = 75
x = 5
-----
since it's small numbers, i tracked it each week as follows:
after week 1
you have 65 and she has 125
after week 2
you have 70 and she has 115
after week 3
you have 75 and she has 105
after week 4
you have 80 and she has 95
after week 5
you have 85 and she has 85
-----